A vertical straight wire carrying an upward 26 A current exerts an attractive force per unit length of 9.0*10^-4 N/m on a second parallel wire 9.0 cm away. What current (magnitude) flows in the second wire?
Answer #1

Force per unit length = F / L = μ0I1I2 / 2πr.

I2 = (F / L)(2πr / μ0I1)

= [(9.0 x 10^-4 N) x ((2π x 0.09 m) / ((4π x 10^-7 Tm/A) x 26 A))

= 15.57 A